### v1

Abstract
Recently, a new general wave phenomenon, namely “the anti-localization of non-stationary linear waves”, has been introduced and discussed. This is zeroing of the propagating component for a non-stationary wave-field near a defect in infinitely long wave-guides. The phenomenon is known to be observed in both continuum and discrete mechanical systems with a defect, provided that the frequency spectrum for the corresponding homogeneous system possesses a stop-band. In this paper, we show that the anti-localization is also quite common for nonlinear systems. To demonstrate this, we numerically solve several non-stationary problems for an infinite $\beta$-FPUT chain with a defect.
1 Introduction
In infinite uniform linear continuum and discrete systems, the perturbations caused by a pulse point load travel as waves [^1] For a discrete system, it is more correct to speak about quasi-waves. along a system from the load application point to infinity. The presence of an inhomogeneity (or a defect) at the load application point can lead to accumulation or decumulation of the wave energy near the defect compared to the uniform case. Apparently, the first study where this fact was demonstrated is the classical paper by Lamb Lamb1900 , where it was shown that embedding a discrete mass-spring system into a continuum medium described by the 1D wave equation leads to the leakage of energy of a mass-spring system. On the other hand, such an embedding leads to the accumulation of energy in the wave-guide near the inhomogeneity, since in a uniform medium, described by the 1D wave equation, the perturbations caused by a point pulse load immediately leave the neighbourhood of the application point, leaving it in the rest [^2] This is due to the structure of the fundamental solution for the 1D wave equation Vladimirov1971 ; Watanabe2015 ..
In uniform linear dispersive systems, the amplitude of perturbations caused by a point pulse load vanishes with time in a continuous manner. Generally, in 1D systems the order of decay for perturbations (e.g., for the particle velocities, the strains, or the displacements if the system is fixed somewhere and does not have rigid modes of motion) is $1/\sqrt{t}$, where $t\to\infty$ is time. This result is well-known; see, e.g., Whitham1999 ; Slepyan1972 , and can be obtained by the asymptotic method of stationary phase applied to a representation of the exact non-stationary solution in the form of a Fourier integral. Again, introducing the inhomogeneity can lead to the accumulation or decumulation of the wave energy nearby.
The effect of the accumulation can be related to the so-called trapped modes ursell1951trapping ; Ind-book-R2E ; Kuznetsov2002 ; Andrianov2013 ; Gavrilov2024nody , which can exist in dispersive systems. In infinite discrete systems (lattices or chains), the analogous phenomenon is known as the impurity-induced localized modes Montroll1955 ; Kivshar1991 ; Kivshar1997 ; Mamalui1999 . The existence of such modes leads to trapping of some portions of the wave energy of a non-stationary process forever near a load application point at an inhomogeneity kaplunov1986torsional ; Teramoto1960 (we call this the localization of non-stationary waves). The whole frequency spectrum of our system can have a stop-band, where complex wave-numbers correspond to real frequencies. In such systems, so-called strong type of localization luongo2001mode ; Luongo1992 is quite common, which is related to a pole in the denominator of the system Green function in the frequency domain lying inside the stop-band. The rest of the energy (or all the wave energy if there is no localization) still travels to infinity as a vanishing propagating component of the wave-field, where perturbations again have the order of decay $1/\sqrt{t}$ (in a 1D system).
Recently Shishkina2023cmat ; Shishkina2023jsv ; Gavrilov2024nody , an opposite general wave phenomenon, which has been called the anti-localization of non-stationary waves, related to the decumulation of the energy near an inhomogeneity, has been introduced. It is observed in both discrete and continuum infinite linear systems, having the dispersion graph with a stop-band [^3] The frequency spectrum of a discrete linear infinite system always has a stop-band Kossevich1999 for high values of the frequency.. The presence of a stop-band leads to the existing of the cut-off frequency to which zero group velocity corresponds. The oscillation corresponding to the cut-off frequency accumulates near a point pulse source hemmer1959dynamic ; slepyan1987energy ; this can cause the phenomenon of the anti-localization. The existence of such a phenomenon is in agreement with some particular results previously obtained kaplunov1986torsional ; Mueller1962 ; Mueller2012 ; Kashiwamura1962 ; Rubin1963 ; hemmer1959dynamic for some systems, where the strong localization can take place. The anti-localization is zeroing of the leading-order term of order $1/\sqrt{t}$ in asymptotic expansions for perturbations at an inhomogeneity. Consequently, the next term of order $1/\sqrt{t^{3}}$ becomes a leading-order term there. Accordingly, the perturbations, which correspond to the propagating component, in an expanding with time neighbourhood of the inhomogeneity, become asymptotically small compared to ones in zones closer to the leading wave-fronts. The amplitude of oscillation in the propagating component has a clear almost zero minimum at the inhomogeneity. Note that the anti-localization can be observed in the absence of the localization as well as co-exist with the localization. In the latter case, after some time, a peak of amplitude for oscillation also becomes clearly recognizable, being observed in the centre of a zone where perturbation amplitude is almost zero.
In the paper, we show that the anti-localization is not a specific peculiarity of linear systems. We demonstrate that it is also common for nonlinear ones. To achieve this, we solve several non-stationary problems for an infinitely long $\beta$-FPUT chain, introduced in the famous study Fermi1955 , with a defect. In the nonlinear case, we have no possibility to investigate the anti-localization in an analytic way using the method of stationary phase. Accordingly, we use a pure numerical approach, considering the previously obtained solutions for the corresponding linear systems as the reference solutions. All numerical calculations presented in the paper have been performed twice using completely independent code and different numerical approaches. The first approach is to use routine scipy.integrate.odeint included in python framework scipy scipy , which is actually based on lsoda from the fortran library odepack. The second approach is related to the self-written code utilizing the semi-implicit Euler method Skeel2020 . The results obtained by both approaches seem almost identical. The choice of a discrete system (not a continuum one) as the system under investigation is caused by the considerations of the simplicity of numerical solving.
In our opinion, the anti-localization of non-stationary waves essentially influences the processes of energy (heat) transfer in linear and nonlinear lattices. In particular, the emergence of cold points at the defects demonstrated in linear and nonlinear cases Gendelman2021 , apparently, can be associated with this phenomenon, though the mathematical formulation used in Gendelman2021 is different. The existence of the anti-localization makes the single-mode approximation very effective Gavrilov2024nody when looking for approximate non-stationary solutions describing non-stationary oscillation in the systems where the strong localization occurs.
The structure of the paper is as follows. In Sect. 2, we discuss the mathematical formulation of the problem in the dimensionless form (the non-dimensionalization is provided in Appendix A). In Sects. 36, several nonlinear problems, where we expect the emergence of the anti-localization, are investigated. In Conclusion (Sect. 7), we discuss the basic results of the paper.
2 The mathematical formulation
Consider an infinite ordered chain of point particles with nearest-neighbour interactions. All particles have an equal mass, whereas one particle has an alternated mass. The particles are labeled by the discrete variable $n\in\mathbb{Z}$, where $\mathbb{Z}$ is the set of all integers. Without loss of generality, we assume that the alternated mass corresponds to $n=0$. The particles are connected by nonlinear springs (i.e., by the bonds) with cubic nonlinearity also labeled by $n\in\mathbb{Z}$. We assume that a spring label equals the label of the preceding particle. All springs have equal stiffness excepting two of them with numbers $n=-1$ and $n=0$, which connect the particle with $n=0$ and its neighbours. Thus, we consider a $\beta$-FPUT chain with a point impurity (or a defect) in the form of a mass-spring inclusion. Everywhere through the paper, excepting Sect. 6, we assume that the inclusion is symmetric [^4] In Sect. 6 we, in particular, consider a linear chain with an inclusion in the form of a nonlinear spring with the linear stiffness being equal to the spring stiffness in the rest of the chain.. The schematic of the system is presented in Fig. 1.
Figure 1: The schematic of the system
The equations of motion in the dimensional form are formulated in Appendix A, where the procedure on non-dimensionalization is presented. In the dimensionless form, we have the following infinite system of differential-difference equations:
$\dot{u}_{n}=v_{n},$ (2.1)
[…]
where
$m_{n}=1+(m-1)\delta_{n},$ (2.5)
$k_{n}=1+(K-1)(\delta_{n+1}+\delta_{n}),$
$\beta_{n}=\beta+(B_{L}-\beta)\delta_{n+1}+(B_{R}-\beta)\delta_{n}.$
Here, overdot denotes the derivative with respect to the dimensionless time $t$; $u_{n}(t)$ is the dimensionless displacement of the particle with a label $n\in\mathbb{Z}$; $\beta$ is the dimensionless cubic stiffness for the bonds outside the defect; $m>0$ is the dimensionless mass of the particle at the defect; $K$ and $B$ are the linear and cubic dimensionless stiffnesses for the bonds at the defect, respectively; $\delta_{n}$ is the Kronecker delta ($1$ if and only if $n=0$, 00 otherwise, $n\in\mathbb{Z}$). The initial conditions are:
$u_{n}(0)=0,\qquad v_{n}(0)=m_{N}^{-1}\delta_{n-N}.$ (2.6)
Initial conditions in the form of Eq. (2.6) correspond to the impulse force $\delta(t)$ applied to the particle with label $N$, see Appendix A. Here $\delta(t)$ is the Dirac delta-function.
Note that we consider the system with a finite nonlinearity, i.e., the dimensionless parameter
$\bar{\beta}=\max_{n}\big{(}|\beta_{n}|\big{)}$ (2.7)
is assumed to be of order one, i.e., not a small quantity. The finite nonlinearity makes linear and nonlinear results to be clearly different for values of time $t$ such that the leading-order term of the perturbation asymptotics calculated for the corresponding linear system by the method of stationary phase eventually becomes close to the numeric results for the linear system.
In the particular case
$m=1,\qquad K=1,\qquad B_{L}=B_{R}=\beta=0$ (2.8)
of an infinite uniform linear chain, the exact solution of the problem defined by Eqs. (2.1)–(2.6) for the particle velocities $v_{n}(t)$ is $V_{n-N}$ schrodinger1914dynamik ; Hamilton1940 ; Havelock1910 ; Muehlich2020 :
$V_{n}=J_{2n}(2t)=J_{2|n|}(2t),$ (2.9)
where $J_{n}(\cdot)$ is the Bessel function of the first kind of order $n\in\mathbb{Z}$.
3 An infinite $\beta$-FPUT chain with an isotopic defect
We assume
$B=B_{L}=B_{R}$ (3.1)
unless the opposite is explicitly indicated. Consider a $\beta$-FPUT chain with an isotopic impurity (a defect)
$m\neq 1,\quad m>0,$ (3.2)
in chain where
$B=\beta,$ (3.3)
$K=1.$ (3.4)
The defect is subjected to a pulse loading at position
$N=0.$ (3.5)
In the linear case $\beta=0$, this problem is considered in studies Rubin1963 ; Shishkina2023cmat . In Shishkina2023cmat an approximate solution of asymptotic nature, describing the wave-field as a whole and valid for large values of time $t$, is obtained. We reproduce here some important formulae from Shishkina2023cmat to explain to the reader what the anti-localization of non-stationary waves is. The exact solution for the particle velocities in the integral form is an inverse Fourier transform:
${v}_{n}=-\frac{\mathrm{i}}{2\pi}\left(\int_{\mathbb{P}}+\int_{\mathbb{S}}\right)\Omega\mathcal{G}_{n}(\Omega)\mathrm{e}^{-\mathrm{i}\Omega t}\,\mathrm{d}\Omega+\mathrm{c.c.}\\ =I_{n}^{\mathrm{pass}}+I_{n}^{\mathrm{stop}}+\mathrm{c.c.}=v_{n}^{\mathrm{pass}}+v_{n}^{\mathrm{stop}},$ (3.6)
where $\mathcal{G}_{n}(\Omega)$ is the corresponding Green function in the frequency domain:
$\mathcal{G}_{n}(\Omega)=-\frac{\mathrm{e}^{\mathrm{i}|n|\operatorname{sign}\Omega\,\arccos\frac{2-\Omega^{2}}{2}}}{(m-1)\Omega^{2}+\mathrm{i}\Omega\sqrt{4-\Omega^{2}}},\ \ \Omega\in\mathbb{P};$ (3.7)
$\mathcal{G}_{n}(\Omega)=\frac{(-1)^{|n|}2^{|n|}}{\Phi^{|n|-1}(\Omega)\big{(}(-m\Omega^{2}+2)\Phi(\Omega)+4\big{)}},\ \ \Omega\in\mathbb{S};$ (3.8)
$\mathrm{c.c.}$ are the complex conjugate terms. Here,
$\Phi(\Omega)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\Omega^{2}-2+|\Omega|\sqrt{\Omega^{2}-4};$ (3.9)
=$\mathbb{P}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}[0,\Omega_{\ast}],\qquad\mathbb{S}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}(\Omega_{\ast},\infty),$ (3.10)
are the pass-band and the stop-band of the whole non-negative frequency band $\Omega\in[0,+\infty]$; $\Omega_{\ast}=2$ is the cut-off frequency which separates the bands.
Remark 3.1
In Shishkina2023cmat mostly the heat, i.e., the kinetic energy, transfer processes are under consideration. Thus, the solution is obtained for the particle velocities, not for the displacements. This is also useful because the displacements involve a non-vanishing non-oscillating component (a rigid body mode of motion).
The integral $I_{n}^{\mathrm{stop}}$ describes, in particular, an impurity-induced non-vanishing localized oscillation, which exists only in the case of a light defect
$10<m<$ 1 . (3.11)
If the last inequality is true, in the interval $(0,\Omega_{\ast})$ there exists a simple root $\Omega_{0}$ of the denominator for $\mathcal{G}^{\mathrm{stop}}$ such that
$\Omega_{0}=\frac{2}{\sqrt{m(2-m)}},$ (3.12)
and, therefore, integral $I^{\mathrm{stop}}$ does not exist in the classical sense. In the latter case, integral $I^{\mathrm{stop}}$ should be considered as the Fourier transform for a generalized function. To regularize this, we can apply Shishkina2023jsv ; Shishkina2023cmat the limit absorption principle. Finally, for $t\to\infty$ at a fixed position $n$ one gets
$v_{n}^{\mathrm{stop}}=I^{\mathrm{stop}}+\mathrm{c.c.}=H(1-m)L_{n}(t)+O(t^{-1}),$ (3.13)
${L}_{n}=\frac{2(m-1)(-1)^{|n|+1}m^{|n|-1}H(1-m)}{(2-m)^{|n|+1}}\cos\Omega_{0}t,$ (3.14)
where $t\to\infty$, see Shishkina2023cmat . Here $L_{n}(t)$ is the localized non-vanishing oscillation, $H(\cdot)$ is the Heaviside step-function.
The phenomenon of the anti-localization is mostly related to the integral $I^{\mathrm{pass}}$, which describes the propagating part of the wave-field. Following to Gavrilov2022ijhmt , we have estimated it in Shishkina2023cmat on a moving at an arbitrary sub-critical speed $w$ point of observation. Taking into account that solution (3.6) clearly is an even function of $n$, put
$|n|=wt,$ (3.15)
$1\displaystyle 0<w<$ 1 (3.16)
in the expression for $I^{\mathrm{pass}}$ defined by Eq. (3.6). The obtained integral can be estimated for $t\to\infty$ using the method of stationary phase. This provides an asymptotic continuum solution:
$v_{n}^{\mathrm{pass}}=I^{\mathrm{pass}}+\mathrm{c.c.}\\ =\frac{{A}(w)}{\sqrt{t}}\cos\Big{(}\phi(w)\,t+\frac{\pi}{4}+\psi(w)\Big{)}+O(t^{-3/2}),$ (3.17)
$\begin{gathered}{A}(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\!\frac{H(1-w)\,w}{\sqrt{\pi}{(1-w^{2})^{1/4}}\big{(}(m-1)^{2}(1-w^{2})+w^{2}\big{)}^{1/2}},\\ \\ \end{gathered}$ (3.21)
$\phi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}2(w\arccos w-\sqrt{1-w^{2}}),$ (3.22)
$\psi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\arctan\frac{(m-1)\sqrt{1-w^{2}}}{w}.$ (3.23)
Note that for $w>1$ there is no stationary point, and the corresponding term of order $1/\sqrt{t}$ is zero. This fact is taken into account by introducing the multiplier $H(1-w)$ in the numerator of the right-hand side of Eq. (3.21). In any system with an isotopic impurity, i.e., in the case (3.2), we have:
${A}(w)=\frac{\sqrt{2}w}{\sqrt{\pi}|m-1|}+O(w^{3}),\quad w\to 0;\qquad A(0)=0;$ (3.24)
and, thus, the amplitude $A(w)$ of the propagating part $v^{\mathrm{pass}}$ for the particle velocities is small in a certain expanding (since $w=|x|/t$) neighbourhood of the impurity. We call this phenomenon the anti-localization of non-stationary waves. Finally, the asymptotic expansion for the right-hand side of Eq. (3.6) at $n=0$ (just at the impurity) has the following form
$v_{0}(t)=H(M-K)L_{0}(t)+\frac{\cos\left(2t-\frac{3\pi}{4}\right)}{2\sqrt{\pi}(m-1)^{2}\,t^{3/2}}\\ +o(t^{-3/2}).$ (3.25)
The first term in the right-hand side of Eq. (3.25) is the contribution from the poles ±$\pm\Omega_{0}$ and describes localized oscillation, which exists if and only if $m<1$. The second term is the anti-localized part of the wave-field at the inclusion expressed as the total contribution from the cut-off frequency $\Omega_{\ast}$ for both integrals $I^{\mathrm{stop}}$ and $I^{\mathrm{pass}}$.
In the absence of the impurity, the anti-localization is not observed:
$A(w)=\frac{1}{\sqrt{\pi{\sqrt{1-w^{2}}}}}=\frac{1}{\sqrt{\pi}}+O(w^{2}),\quad w\to 0;\\ \qquad A(0)\neq 0.$ (3.26)
Though the last formula can be formally derived from Eq. (3.21), this way is incorrect since Eq. (3.21) is obtained assuming Eq. (3.2). The correct way to obtain formula (3.26) is to use the exact solution (2.9) Gavrilov2022ijhmt . Thus, the anti-localization in the linear system is an asymptotic weakening of the propagating component of the wave-field. The order of decay for the vanishing component is $1/\sqrt{t}$ in the uniform system and $1/\sqrt{t^{3}}$ in the system with a defect (near the defect).
Now let us consider how a nonlinearity influences the anti-localization. To do this, following to Shishkina2023cmat , we proceed with numerical integration of $12n_{0}+$ 1 ODE (2.1), (2.4) with $n\in-n_{0}\dots n_{0}$, $n_{0}>\lambda t$, $\lambda=\operatorname{const}>1$. At first, consider the case of a heavy defect, where the impurity-induced localized mode does not exist for the corresponding linear system with $\beta_{n}=0$, and, hence, $v_{n}^{\mathrm{stop}}=0$.
In Fig. 2, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time. For reference, the wave-field in the corresponding linear system with $\beta_{n}=0$ given by Eqs. (3.6), (3.13)–(3.23), the wave-field in the uniform linear chain given by Eq. (2.9), and the numerically found wave-field in the corresponding uniform $\beta$-FPUT chain are also displayed. One can compare Fig. 2 with Fig. 2a in Shishkina2023cmat , where it is demonstrated that the numeric results for the linear system are in excellent agreement with asymptotic description Eqs. (3.6), (3.13)–(3.23), everywhere excepting the leading wave-fronts $n\simeq t$. One can see that nonlinearity essentially affects the wave-fields near the leading fronts $n\simeq t$, see Li2010 . However, in a central zone far from the leading fronts, the presence of nonlinearity does not alter the qualitative properties of the wave-fields essentially. Zeroing of the oscillation amplitude near $n=0$, i.e., the anti-localization, is observed in both linear and nonlinear systems with the defect. On the other hand, the anti-localization is not observed in both uniform systems under investigation.
In Fig. 3a the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the heavy defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (3.25) is also displayed. According to Fig. 3a, the order of decay for oscillation in the nonlinear system can be estimated the same as it is observed in the corresponding linear system, i.e., $t^{-3/2}$, see Eq. (3.25). The essential peculiarity of the oscillation in the nonlinear system with the defect, which can be observed in Fig. 3, is the oscillation asymmetry. Indeed, one can see that there exists a vanishing (apparently with the same power law $t^{-3/2}$) non-negative component of the particle velocity $v_{0}(t)$. Such a component does not exist in the corresponding linear system.
In Fig. 3b the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system without the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (2.9), is also displayed. The order of decay for oscillation in the nonlinear system can be estimated as $t^{-1/2}$ as it is observed according to Eqs. (3.21), (3.22), (3.26) in the uniform linear system. Thus, there is no anti-localization in both nonlinear and linear systems without a defect.
Finally, consider the case of a light defect $10<m<$ 1. As it is shown in Shishkina2023cmat , in the corresponding linear system, the localized component ${v}_{n}^{\mathrm{stop}}$ co-exists with the anti-localized one ${v}_{n}^{\mathrm{pass}}$. We also know that adding nonlinearity does not destroy an impurity-induced localized mode in the case of the chain Kivshar1991 . In Fig. 4, where the particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time, one can see that in the nonlinear system we also observe a similar wave pattern, which corresponds to the superposition of localized and anti-localized components. In Fig. 5, the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (3.25) is also displayed. In both cases (the nonlinear and the linear ones), a non-vanishing impurity induced localized oscillation is observed.
Qualitatively similar results can be exhibited for various values of the system parameters, such as for a hard $\beta$-FPUT chain with $\beta>0$ as well as for a soft one with $\beta<0$.
Figure 2: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a heavy isotopic defect. The anti-localization, i.e., zeroing of the perturbation amplitude near $n=0$, is clearly observed for both linear and nonlinear systems with the defect and not observed for systems without the defect
Figure 3: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solutions for a system with a heavy isotopic defect; (b) the solutions for the corresponding uniform system. In both sub-plots, the order of decay for the nonlinear system can be estimated as the same as it is observed in the corresponding linear system
Figure 4: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a light isotopic defect. The anti-localization in systems with the defect co-exists with impurity-induced localized mode. The latter forms a peak of the amplitude near $n=0$
Figure 5: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solution for a system with a light isotopic defect; (b) the solution for the corresponding uniform system. For both linear and nonlinear systems, a non-vanishing localized oscillation is observed
4 An infinite $\beta$-FPUT chain with a mass-spring inclusion
As we can see from the previous Sect. 3, the anti-localization exists in linear and nonlinear systems with an isotopic defect. The anti-localization is observed for both cases $m>1$ and $m<1$, though, in the latter case, the anti-localization co-exists with the localization. The boundary in the system parameter space between these two cases corresponds to the uniform system, where there is no anti-localization. In previous study Shishkina2023jsv , we have shown that besides the uniform system, the anti-localization disappears in linear non-uniform systems with parameters, which correspond to such a boundary separating domains in the system parameter space where the localization is observed and not observed. The linear uniform system is just a particular case of such a system with the boundary parameter values. In this section, we want to check if the anti-localization disappears for non-uniform nonlinear systems with such boundary values of their parameters. To achieve this, we consider the chain with parameters satisfying Eq. (3.3), equipped with a mass-spring inclusion for which Eq. (3.2) and
$K\neq 1$ (4.1)
are fulfilled. The defect is subjected to a pulse loading at the position defined by Eq. (3.5). In the corresponding linear system with $B=\beta=0$ the impurity-induced localization is observed if and only if the following inequality is fulfilled Montroll1955 ; Yu2019 ; Shishkina2023asm :
$K>K_{0},$ (4.2)
$K_{0}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\frac{2m}{1+m}.$ (4.3)
Inequality (4.2) defines a domain in the problem parameter space, which we call the localization domain. Now consider a nonlinear system with a fixed value $B=\beta$, and fixed $m$ as we have taken in Sect. 3. Now we try to determine the boundary value of the parameter $K$, which separates nonlinear systems where the impurity-induced localization is observed or not observed. The numeric experiment shows that such a value is very close to the value $K=K_{0}$ defined by Eq. (4.3) for the linear systems, see Figs. 68.
Figure 6: The particle velocity $v_{0}(t)$ versus time in the case of the mass-spring defect. Parameters of defect satisfy equation $K=K_{0}$. The order of decay for all three systems can be estimated as the same
Figure 7: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in the case of the mass-spring defect. Parameters of defect satisfy equation $K=K_{0}$. The anti-localization is absent in all three cases (the minimum at $n=0$ observed for the linear systems is not an amplitude minimum and related to the value of the phase $\phi t$)
Figure 8: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain with a mass-spring defect. The orders of decay are clearly different in all three cases.
In Fig. 6 the numerical solution $v_{0}(t)$ versus the time $t$ for $\beta$-FPUT chain with a mass-spring defect is presented. The value of the linear spring stiffness is taken equal to the boundary value $K_{0}$. For reference, the same particle velocity in the corresponding linear system with the same defect, as well as the exact solution (2.9) for the uniform linear chain, are also presented. [^5] Unfortunately, we do not have here the reference asymptotic non-stationary solution for the corresponding linear system. The order of decay for oscillation in both nonlinear systems with and without the defect can be estimated as $1/\sqrt{t}$ as it is observed according to Eqs. (3.17), (3.26) in the uniform linear chain. In Fig. 7, the particle velocities $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ are shown at a fixed value of time for the same three systems as in Fig. 6. The absence of the anti-localization in the system with the defect for which $K=K_{0}$ is clearly recognizable.
In Fig. 8 the same numerical solution $v_{0}(t)$ for $\beta$-FPUT chain with a mass-spring defect with $K=K_{0}$, which is already shown in Fig. 6, is compared with analogous solutions with $K=0.9K_{0}$ and $K=1.1K_{0}$. We can see that the solution corresponding to $K=0.9K_{0}$ vanishes faster than one taken for $K=K_{0}$. The corresponding order of decay is expected to be $1/\sqrt{t^{3}}$, see Sect. 3. On the other hand, the solution corresponding to $K=1.1K_{0}$ is a non-vanishing one, i.e., an impurity induced localized oscillation is observed.
Thus, we have shown that the anti-localization is not observed in both linear and nonlinear systems with a mass-spring defect if the system parameters correspond to the boundary of the localization domain in the problem parameter space. On the other hand, the anti-localization is observed if the system parameters lie outside this boundary.
5 A semi-infinite $\beta$-FPUT chain
Assuming that $K=0$, $B=0$, we split an infinite chain into two uncoupled semi-infinite chains. Thus, a semi-infinite chain loaded at the end $n=N=1$ is a particular case of a chain with a defect. The exact solution is Liazhkov2025
$v_{n}=J_{2|n-1|}(2t)+J_{2n}(2t).$ (5.1)
The anti-localization in such a linear chain ($\beta=0$) was demonstrated in study Gavrilov2023DD , where the asymptotic solution for $v_{n}$ in the form of the right-hand side of Eq. (3.17) wherein
${A}(w)=\frac{2w}{\sqrt{\pi}\sqrt[4]{1-w^{2}}},$ (5.2)
$\psi=\arctan\frac{\sqrt{1-w^{2}}}{w},$ (5.3)
$w=\frac{|n-1|}{t}$ (5.4)
was obtained. The expression for $\phi$ coincides with (3.22). Clearly, $A(0)=0$, which implies the anti-localization effect. For the end particle, one has Gavrilov2023DD :
$v_{1}=\frac{\sin(2t-\frac{\pi}{4})}{\sqrt{\pi}t^{3/2}}+o(t^{-3/2}).$ (5.5)
This effect was discovered for the first time, as far as we know now [^6] We are grateful to S.D. Liazhkov, much earlier in Jackson1978 and explained by specific peculiarities of linear systems. Let us check if the anti-localization exists in a semi-infinite $\beta$-FPUT chain.
The numerical solution, which corresponds to a nonlinear chain with $\beta\neq 0$ is presented in Fig. 9. For reference, the wave-field in the corresponding linear semi-infinite chain given by the exact solution (5.1) is also displayed. In Fig. 9a the particle velocities $v_{n}$ versus the spatial variable $n$ are presented. In Fig. 9b the particle velocity of the end particle $v_{1}(t)$ versus time is shown. One can see that the anti-localization is clearly visible for both linear and nonlinear systems.
Remark 5.1
In Fig. 9a one can also observe a soliton-like wave propagating faster than the leading wave-front in the linear system. The solitons in a $\beta$-FPUT chain are discussed, e.g., in Vainchtein2022 .
Figure 9: The wave-field in the $\beta$-FPUT and linear chains loaded at the free end: (a) particle velocities $v_{n}$ versus the spatial variable $n$; (b) particle velocity of the end particle $v_{1}$ versus time. The anti-localization is clearly observed
6 A nonlinear spring in an infinite linear chain
In this section, we want to show that introducing the local nonlinearity into a linear system can also lead to a phenomenon similar to the anti-localization. We consider a uniform linear chain, i.e.,
$m=1,\quad K=1,\quad\beta=0.$ (6.1)
At first, consider the source at $N=0$ at the inclusion in the form of a hard nonlinear spring:
$B_{L}=0,\quad B_{R}>0.$ (6.2)
In Fig. 10 we present the spatial distribution of the particle velocities in the corresponding numeric experiment with pulse loading applied at the particle with number $n=N=0$ just to the left from the nonlinear spring. Sub-plots (a) & (b) correspond to the choices $B_{R}=4.0$ and $B_{R}=10.0$, respectively. For reference, the particle velocities in the corresponding linear system with $B_{L}=B_{R}=0$ given by Eq. (2.9) are also displayed in each sub-plot. One can see that the wave pattern observed in Fig. 10b is a bit similar to the one presented in Fig. 4, where the anti-localization co-exists with the impurity-induced localized oscillation. The clear difference is an expected asymmetry of the wave patterns in Fig. 10. In Fig. 10a we also observe weakening of the propagating component near the defect but not zeroing. Increasing the coefficient $B_{R}$ one can see that the amplitude of the propagating component near the defect decreases. Thus, we observe a weakening of the propagating component but, apparently, not an asymptotic one. The observed phenomenon can be called the incomplete anti-localization. Here we cannot estimate, even qualitatively, the order of decay for the anti-localized component due to the presence of the localized mode.
Figure 10: The particle velocity $v_{n}$ versus the spatial variable $n$ in an infinite linear chain with and without a nonlinear spring. The greater $B_{R}$, the tendency to the anti-localization is more clear
A possibility to eliminate the localized oscillation is to consider a pulse loading applied at some distance $N>0$ outside the defect. For such a case, in study Gavrilov2024 , we have shown that in the linear chain the anti-localization is observed in some one-sided neighbourhood just behind the defect, which expands with time. The amplitude of the localized oscillation in the linear system exponentially decreases with $N$; therefore this oscillation is not observable for large enough $N$. The order of decay for the particle velocities in such a shadow zone is the same as it is observed when the loading is applied at a defect, i.e., $t^{-3/2}$. In Fig. 11, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ in an infinite linear chain, with an inclusion in the form of one strongly nonlinear spring, is shown for a fixed value of time. Here, we take an extremely large value $B_{R}=1000$ to obtain a clearly visible effect. For reference, the wave-field in the uniform linear chain given by Eq. (2.9), is also displayed. One can compare Fig. 11 with Fig. 1 in Shishkina2023cmat , where the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for the analogous problem for the linear chain with an isotopic defect. The plots are very similar, and the shadow zones behind the defect are clearly recognizable in both cases.
Figure 11: The particle velocity $v_{n}$ versus the spatial variable $n$ in an infinite linear chain with a hard, strongly nonlinear spring. The source position $n=N$ is indicated by the vertical magenta solid line. One can clearly see a shadow behind the defect at $n=0$
In Fig. 12 the numerically found solution for the particle velocity $v_{-1}(t)$ versus the time $t$ is shown for the same system. The choice $n=-1$ corresponds to the first particle behind the defect, which is not connected to the defective spring. One can see that the estimate for the order of the decay should be taken as $t^{-1/2}$, not $t^{-3/2}$ as usual when the (complete) anti-localization is observed. Thus, we really observe a weakening of the oscillation behind the defect, but not an asymptotic one. The leading order term of the large time asymptotics for the oscillation amplitude is relatively small, but not zero.
Figure 12: The particle velocity $v_{-1}$ versus time in an infinite linear chain with a nonlinear spring. The order of decay can be estimated as $1/\sqrt{t}$
7 Conclusion
In the paper, we have presented the numerical results for investigation of several nonlinear problems concerning the quasi-waves propagation in an infinitely long $\beta$-FPUT chain with a defect. Though we have demonstrated the results, which correspond to a specific choice of the problem parameters, in our work, we have considered a wide band of them. This allows us to claim that the phenomenon of anti-localization of non-stationary waves is not somehow related to the linearity of the system under consideration. Anti-localization is natural for both linear and nonlinear systems with defects.
Depending on the parameters of the system, we have the following alternatives:
  1. 1.
    An anti-localized propagating component of a wave-field may co-exist with an impurity-induced localized wave; see, e.g., Fig. 4, which corresponds to a light isotopic defect. The localized mode forms a peak of the amplitude near a defect, but this peak is observed inside a zone, where the oscillation amplitude is zero.
  2. 2.
    An anti-localized propagating component can be observed near a defect in the absence of impurity-induced localized modes; see, e.g., Fig. 2, which corresponds to a heavy isotopic defect.
  3. 3.
    The anti-localization is not observed for systems with parameters at the boundary of the domain in the problem parameter space where the wave localization occurs; see, e.g., Fig. 7, which corresponds to a mass-spring defect. Thus, we have shown that the absence of the anti-localization in a uniform system is not related to the uniformity itself. A uniform system is just the particular case of a system with such boundary parameters.
This was shown previously in Shishkina2023jsv for a continuum linear system in an analytic way. In this paper, we have demonstrated numerically an analogous behaviour of nonlinear systems with finite nonlinearity.
In Sect. 6, we have demonstrated that embedding the point strong nonlinearities into a linear chain can also lead to a weakening of the propagating component of the wave-field in a neighbourhood of such a defect. Though this effect is quite similar to one which we call the anti-localization, in the latter case, we observe weakening but not zeroing of the leading-order term. This can be called “the incomplete anti-localization”.
Note that introducing nonlinearity into a linear system possessing a localized mode with the corresponding frequency inside a stop-band, in some cases, can lead to destroying the mode. This happens at least in the case when the stop-band correspond to smaller values of frequency compared with the pass-band kaplunov2008example ; indeitsev2004localization due to the emergence of super-linears, whose frequencies are multiples of the localized mode frequency. These frequencies can leave the stop-band yielding the leakage of energy of the localized mode. In this paper, we consider systems with a stop-band corresponding to higher values of the frequency compared with a pass-band. This situation is quite usual for discrete systems with zero on-site potential. The anti-localization is the phenomenon, which is closely related to the localization; thus, in nonlinear systems where a stop-band corresponds to small values of the frequency (this is typical for the continuous case) the anti-localization requires an additional investigation.
Acknowledgements
The authors are grateful to E.F. Grekova, S.D. Liazhkov, Yu.A. Mochalova, A.V. Porubov for useful and stimulating discussions.
Funding
The research is supported by the Ministry of Science and Higher Education of the Russian Federation (project 124040800009-8).
Appendix A Non-dimensionalization
In the dimensional form, the equations of motion are:
$m_{\infty}\frac{\mathrm{d}^{2}{\tilde{u}}_{n}}{\mathrm{d}\tilde{t}^{2}}-K_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})-(\tilde{u}_{n}-\tilde{u}_{n-1})\big{)}\\ -B_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})^{3}-(\tilde{u}_{n}-\tilde{u}_{n-1})^{3}\big{)}\\ =\Big{(}(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})-(\tilde{u}_{0}-\tilde{u}_{-1})\big{)}\\ +\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\big{)}\Big{)}\delta_{n}\\ -(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})\delta_{n-1}-(\tilde{u}_{0}-\tilde{u}_{-1})\delta_{n+1}\big{)}\\ -\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}\delta_{n-1}\\ -(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\delta_{n+1}\big{)}\\ -(\tilde{m}-m_{\infty})\frac{\mathrm{d}^{2}{\tilde{u}}_{0}}{\mathrm{d}\tilde{t}^{2}}\delta_{n}+\tilde{p}\delta_{n-N}\delta(\tilde{t}).$ (A.1)
Here, $n\in\mathbb{Z}$; $\tilde{t}$ is the time, $\tilde{u}_{n}(\tilde{t})$ is the displacement of the particle with a number $n$; $m_{\infty}$ is the mass for particles outside the defect; $K_{\infty}$, $B_{\infty}$ are the linear and cubic stiffnesses for the bonds outside the defect; $\tilde{m}$ is the mass of the particle at the defect, $\tilde{K}$ is the linear stiffness for the bonds at the defect; $\tilde{B}_{L}$ and $\tilde{B}_{R}$ are the cubic stiffnesses for the bonds to the left and to the right of the defect. Since the right-hand side of Eq. (A.1) involves the term proportional to the Dirac delta-function $\delta(t)$, we formulate initial conditions in the form, which is conventional in the theory of generalized functions (distributions):
$\tilde{u}_{n}\big{|}_{t<0}\equiv 0.$ (A.2)
Put
$\begin{gathered}\omega=\sqrt{\frac{K_{\infty}}{m_{\infty}}},\quad t=\omega\tilde{t},\quad u_{n}=\frac{\tilde{u}_{n}K_{\infty}}{\tilde{p}\omega},\\ K=\frac{\tilde{K}}{K_{\infty}},\quad m=\frac{\tilde{m}}{m_{\infty}},\\ B_{L}=\frac{\tilde{B}_{L}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad B_{R}=\frac{\tilde{B}_{R}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad\beta=\frac{B_{\infty}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}}.\end{gathered}$ (A.3)
Using dimensionless quantities (A.3), Eq. (A.1) may be rewritten in the following dimensionless form:
$\ddot{u}_{n}-(u_{n+1}-2u_{n}+u_{n-1})\\ -\beta\big{(}(u_{n+1}-u_{n})^{3}-(u_{n}-u_{n-1})^{3}\big{)}\\ =\big{(}(K-1)\big{(}(u_{1}-u_{0})-(u_{0}-u_{-1})\big{)}\\ +\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\big{)}\big{)}\delta_{n}\\ -(K-1)\big{(}(u_{1}-u_{0})\delta_{n-1}-(u_{0}-u_{-1})\delta_{n+1}\big{)}\\ -\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}\delta_{n-1}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\delta_{n+1}\big{)}\\ -(m-1)\ddot{u}_{0}\delta_{n}+\delta_{n-N}\,\delta(t),$ (A.4)
which is equivalent to Eqs. (2.1)–(2.4) if notation (2.5) is accepted. The initial conditions (A.2) can be reformulated in the classical form (2.6).