### v2

# Anti-localization of non-stationary quasi-waves in a strongly nonlinear $\beta$-FPUT chain

Abstract
Recently, a new general wave phenomenon, namely ‘‘the anti-localization of non-stationary linear waves’’, has been introduced and discussed (Shishkina et al., J. Sound. Vib. 553, 2023, 117673). This is zeroing of the propagating component for a non-stationary wave-field near a defect in infinitely long wave-guides. The phenomenon is known to be observed in both continuum and discrete mechanical systems with a defect, provided that the frequency spectrum for the corresponding homogeneous system possesses a stop-band. In this paper, we show that the anti-localization is also quite common for nonlinear systems. To demonstrate this, we numerically solve several non-stationary problems for an infinite strongly nonlinear $\beta$-FPUT chain with a defect. In our opinion, the anti-localization essentially influences the processes of heat transfer in linear and nonlinear lattices.
I Introduction
In infinitely long uniform linear continuum and discrete systems, the perturbations caused by a pulse point load travel as waves  [^1] For a discrete system, it is more correct to speak about quasi-waves. along a system from the load application point to infinity. The presence of an inhomogeneity (or a defect) at the load application point can lead to accumulation or decumulation of the wave energy near the defect compared to the uniform case. Apparently, the first study where this fact was demonstrated is the classical paper by Lamb Lamb (1900), where it was shown that attaching a discrete mass-spring system into a continuum medium described by the 1D wave equation leads to the leakage of energy of a mass-spring system. On the other hand, such an attaching leads to the accumulation of energy in the wave-guide near the inhomogeneity, since in a uniform medium, described by the 1D wave equation, the perturbations caused by a point pulse load immediately leave the neighborhood of the application point, leaving it in the rest  [^2] This is due to the structure of the fundamental solution for the 1D wave equation Vladimirov (1971); Watanabe (2015)..
In uniform linear dispersive systems, the amplitude of perturbations caused by a point pulse load vanishes with time not immediately but in a continuous manner. Generally, in 1D systems the order of decay for perturbations (e.g., for the particle velocities, the strains, or the displacements if the system is fixed somewhere and does not have rigid modes of motion) is $1/\sqrt{t}$, where $t\to\infty$ is time. This result is well-known; see, e.g., Whitham (1999); Slepyan (1972), and can be obtained by the asymptotic method of stationary phase applied to a representation of the exact non-stationary solution in the form of a Fourier integral. Again, introducing the inhomogeneity can lead to the accumulation or decumulation of the wave energy nearby.
The effect of the accumulation can be related to the so-called trapped modes Ursell (1951); Indeitsev et al. (2007); Kuznetsov et al. (2002); Andrianov et al. (2013); Gavrilov et al. (2024), which can exist in dispersive systems. In infinite discrete systems (lattices or chains), the analogous phenomenon is known as the impurity-induced localized modes Montroll and Potts (1955); Kivshar (1991); Kivshar et al. (1997); Mamalui et al. (1999). The existence of such modes leads to trapping of some portions of the wave energy of a non-stationary process forever near a load application point at an inhomogeneity Kaplunov (1986); Teramoto and Takeno (1960) (we call this the localization of non-stationary waves). The whole frequency spectrum of our system can have a stop-band, where complex wave-numbers correspond to real frequencies. In such systems, so-called strong type of localization Luongo (2001, 1992) is quite common, which is related to a pole in the denominator of the system Green function in the frequency domain lying inside the stop-band. The rest of the energy (or all the wave energy if there is no localization) still travels to infinity as a vanishing propagating component of the wave-field, where perturbations again have the order of decay $1/\sqrt{t}$ (in a 1D system).
Recently Shishkina and Gavrilov (2023a); Shishkina et al. (2023); Gavrilov et al. (2024), an opposite general wave phenomenon, which has been called the anti-localization of non-stationary waves, related to the decumulation of the energy near an inhomogeneity, has been introduced. It is observed in both discrete and continuum infinite linear systems, having the dispersion graph with a stop-band  [^3] The frequency spectrum of a discrete linear infinite system always has a stop-band Kossevich (1999) for high values of the frequency.. The presence of a stop-band leads to the existing of the cut-off frequency to which zero group velocity corresponds. The oscillation corresponding to the cut-off frequency accumulates near a point pulse source Hemmer (1959); Slepyan and Tsareva (1987); this can cause the phenomenon of the anti-localization. The existence of such a phenomenon is in agreement with some particular results previously obtained Kaplunov (1986); Hemmer (1959); Müller (1962); Müller and Weiss (2012); Kashiwamura (1962); Rubin (1963); Jackson (1978) for some systems, where the strong localization can take place. The anti-localization is zeroing of the leading-order term of order $1/\sqrt{t}$ in asymptotic expansions for perturbations at an inhomogeneity. Consequently, the next term of order $1/\sqrt{t^{3}}$ becomes a leading-order term there. Accordingly, the propagating perturbations, in an expanding with time neighborhood of the inhomogeneity, become asymptotically small compared to ones in zones closer to the leading wave-fronts. The amplitude of oscillation in the propagating component has a clear almost zero minimum at the inhomogeneity. The anti-localization can be observed in the absence of the localization as well as co-exist with the localization. In the latter case, after some time, a peak of the amplitude for oscillation also becomes clearly recognizable, being observed in the center of a zone where the perturbation amplitude is almost zero. Note that the term ‘‘weak anti-localization’’ is commonly known in modern physics Bergmann (1982). The anti-localization of non-stationary waves and the weak anti-localization are different phenomena. The weak anti-localization is related to the weak localization, whereas the anti-localization of non-stationary waves is observed in systems where the strong localization can be possible Shishkina et al. (2023).
In the paper, we show that the anti-localization is not a specific peculiarity of linear systems. We demonstrate that it is also common for nonlinear ones. To achieve this, we solve several non-stationary problems for an infinitely long $\beta$-FPUT chain, introduced in the famous study Fermi et al. (1955), with a defect. In the strongly nonlinear case, we have no possibility to investigate the anti-localization in an analytic way using the method of stationary phase. Accordingly, we use a pure numerical approach, considering the previously obtained solutions for the corresponding linear systems as the reference ones. All numerical calculations presented in the paper have been performed twice using completely independent code and different numerical approaches. The first approach is to use routine scipy.integrate.odeint included in python framework scipy Virtanen et al. (2020), which is actually based on lsoda from the fortran library odepack. The second approach is related to the self-written code utilizing the semi-implicit Euler method Skeel and Cieśliński (2020). The results obtained by both approaches seem almost identical. The choice of a discrete system (not a continuum one) as the system under investigation is caused by the considerations of the simplicity of numerical solving.
In our opinion, the anti-localization essentially influences the processes of heat transfer in linear and nonlinear lattices. In particular, the anti-localization can explain the occurrence of the Kapitza thermal resistance Gavrilov and Shishkina (2024); Gendelman and Paul (2021). The emergence of cold points at the defects demonstrated in linear and nonlinear cases Gendelman and Paul (2021), apparently, can be associated with this phenomenon, though the mathematical formulation used in Gendelman and Paul (2021) is different. The anti-localization can prevent the direct measurement of nonequilibrium temperature of a lattice as discussed in Hatano and Jou (2003), since the attaching the thermometer to the lattice leads to the distortion of the temperature field due to the anti-localization. The existence of the anti-localization makes a single-mode ansatz very effective Gavrilov et al. (2024) when looking for approximate non-stationary solutions describing non-stationary oscillation in the systems where the strong localization occurs.
The structure of the paper is as follows. In Sect. II, we discuss the mathematical formulation of the problem in the dimensionless form. In Sects. IIIVI, several nonlinear problems, where we expect the emergence of the anti-localization, are investigated. In Conclusion (Sect. VII), we discuss the basic results of the paper. In Appendix A, we provide the procedure of the non-dimensionalization for the problem. In Appendix B, we reproduce some important formulas from Shishkina and Gavrilov (2023a) to explain to the reader what the anti-localization of non-stationary waves is in the linear case. In Appendix C, we provide some reference formulas for the semi-infinite linear chain.
II The mathematical formulation
Consider an infinite ordered chain of point particles with nearest-neighbor interactions. All particles have an equal mass, whereas one particle has an alternated mass. The particles are labeled by the discrete variable $n\in\mathbb{Z}$, where $\mathbb{Z}$ is the set of all integers. Without loss of generality, we assume that the alternated mass corresponds to $n=0$. The particles are connected by nonlinear springs (i.e., by the bonds) with cubic nonlinearity also labeled by $n\in\mathbb{Z}$. We assume that a spring label equals the label of the preceding particle. All springs have equal stiffness excepting two of them with numbers n=1n=-1 and $n=0$, which connect the particle with $n=0$ and its neighbors. Thus, we consider a $\beta$-FPUT chain with a point impurity (or a defect) in the form of a mass-spring inclusion. Everywhere through the paper, excepting Sect. VI, we assume that the inclusion is symmetric  [^4] In Sect. VI we, in particular, consider a linear chain with an inclusion in the form of a nonlinear spring with the linear stiffness being equal to the spring stiffness in the rest of the chain.. The schematic of the system is presented in Fig. 1.
Figure 1: The schematic of the system
The equations of motion in the dimensional form are formulated in Appendix A, where the procedure on non-dimensionalization is presented. In the dimensionless form, we have the following infinite system of differential-difference equations:
$\dot{u}_{n}=v_{n},$ (1)
[…]
where
$m_{n}=1+(m-1)\delta_{n},$ (5)
$k_{n}=1+(K-1)(\delta_{n+1}+\delta_{n}),$
$\beta_{n}=\beta+(B_{L}-\beta)\delta_{n+1}+(B_{R}-\beta)\delta_{n}.$
Here, overdot denotes the derivative with respect to the dimensionless time $t$; $u_{n}(t)$ and $v_{n}(t)$ are the dimensionless displacement and particle velocity, respectively, of the particle with a label $n\in\mathbb{Z}$; $\beta$ is the dimensionless cubic stiffness for the bonds outside the defect; $m>0$ is the dimensionless mass of the particle at the defect; $K$ and $B$ are the linear and cubic dimensionless stiffnesses, respectively, for the bonds at the defect; $\delta_{n}$ is the Kronecker delta ($1$ if and only if $n=0$, 0 otherwise, $n\in\mathbb{Z}$).
The initial conditions are:
$u_{n}(0)=0,\qquad v_{n}(0)=m_{N}^{-1}\delta_{n-N}.$ (6)
Initial conditions in the form of Eq. (6) correspond to the pulse loading $\delta(t)$ applied to the particle with label $N$. Here $\delta(t)$ is the Dirac delta-function.
Note that we consider the strongly nonlinear system, i.e., the dimensionless parameter
$\bar{\beta}=\max_{n}\big{(}|\beta_{n}|\big{)}$ (7)
is assumed to be of order one, i.e., not a small quantity. Thus, we expect that the clear difference between a nonlinear solution and the corresponding linear one can be observed for finite values of time $t$.
In the particular case
$m=1,\qquad K=1,\qquad B_{L}=B_{R}=\beta=0$ (8)
of an infinite uniform linear chain, the exact solution of the problem defined by Eqs. (1)–(6) for the particle velocities $v_{n}(t)$ is $V_{n-N}$ Schrödinger (1914); Hamilton (1940); Havelock (1910); Mühlich et al. (2020):
$V_{n}=J_{2n}(2t)=J_{2|n|}(2t),$ (9)
where $J_{n}(\cdot)$ is the Bessel function of the first kind of order $n\in\mathbb{Z}$.
III An infinite $\beta$-FPUT chain with an isotopic defect
We assume
$B=B_{L}=B_{R}$ (10)
unless the opposite is explicitly indicated. Consider a $\beta$-FPUT chain with an isotopic impurity (a defect)
$m\neq 1,\quad m>0,$ (11)
in chain where
$B=\beta,$ (12)
$K=1.$ (13)
The defect is subjected to a pulse loading at position
$N=0.$ (14)
In the linear case $\beta=0$, this problem is considered in studies Rubin (1963); Shishkina and Gavrilov (2023a). In Shishkina and Gavrilov (2023a) an approximate solution of asymptotic nature, describing the wave-field as a whole and valid for large values of time $t$, is obtained. We reproduce in Appendix B some important formulas from Shishkina and Gavrilov (2023a) to explain to the reader what the anti-localization of non-stationary linear waves is.
Now let us consider how a nonlinearity influences the anti-localization. To do this, following to Shishkina and Gavrilov (2023a), we proceed with numerical integration of $2n_{0}+1$ ODE (1), (4) with $n\in-n_{0}\dots n_{0}$, $n_{0}>\lambda t$, $\lambda=\operatorname{const}>1$. At first, consider the case of a heavy defect, where the impurity-induced localized mode does not exist for the corresponding linear system with $\beta_{n}=0$, and, hence, $v_{n}^{\mathrm{stop}}=0$.
In Fig. 2, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time. For reference, the wave-field in the corresponding linear system with $\beta_{n}=0$ given by Eqs. (24), (31)–(41), the wave-field in the uniform linear chain given by Eq. (9), and the numerically found wave-field in the corresponding uniform $\beta$-FPUT chain are also displayed. One can compare Fig. 2 with Fig. 2a in Shishkina and Gavrilov (2023a), where it is demonstrated that the numeric results for the linear system are in excellent agreement with asymptotic description Eqs. (24), (31)–(41), everywhere excepting the leading wave-fronts $n\simeq t$. The nonlinearity essentially affects the wave-fields near the leading fronts $n\simeq t$, see Li et al. (2010). In a central zone far from the leading fronts, we observe that the nonlinearity quantitatively alters the solution: the amplitude of oscillation in the nonlinear case is approximately a half of the amplitude in the linear case. Zeroing of the oscillation amplitude near $n=0$, i.e., the anti-localization, is observed in both linear and nonlinear systems with the defect. On the other hand, the anti-localization is not observed in both uniform systems under investigation.
In Fig. 3a the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the heavy defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$, i.e., the asymptotics given by Eq. (43) and the numerical solution, are also displayed. One can see that in the linear system the numerical solution very often converges to the asymptotics. On the other hand, we clearly deal with an essentially nonlinear regime, since the solutions in linear and nonlinear cases are not close to each other for finite times. According to Fig. 3a, the order of decay for oscillation in the nonlinear system can be estimated the same as it is observed in the corresponding linear system, i.e., $t^{-3/2}$, see Eq. (43). The essential peculiarity of the oscillation in the nonlinear system with the defect, which can be observed in Fig. 3, is the oscillation asymmetry. Indeed, one can see that there exists a vanishing (apparently with the same power law $t^{-3/2}$) non-negative component of the particle velocity $v_{0}(t)$. Such a component does not exist in the corresponding linear system Shishkina and Gavrilov (2023a).
In Fig. 3b the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system without the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (9), is also displayed. The order of decay for oscillation in the nonlinear system can be estimated as $t^{-1/2}$ as it is observed according to Eqs. (39), (40), (44) in the uniform linear system. Thus, there is no anti-localization in both nonlinear and linear systems without a defect.
Finally, consider the case of a light defect $0<m<1$. As it is shown in Shishkina and Gavrilov (2023a), in the corresponding linear system, the localized component ${v}_{n}^{\mathrm{stop}}$ co-exists with the anti-localized one ${v}_{n}^{\mathrm{pass}}$. We also know that adding nonlinearity does not destroy an impurity-induced localized mode in the case of the chain Kivshar (1991). In Fig. 4, where the particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time, one can see that in the nonlinear system we also observe a similar wave pattern, which corresponds to the superposition of localized and anti-localized components. In Fig. 5, the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (43) is also displayed. In both cases (the nonlinear and the linear ones), a non-vanishing impurity induced localized oscillation is observed. The frequency shift observed for the nonlinear system confirms the nonlinear character of the localized oscillation.
Qualitatively similar results can be exhibited for various values of the system parameters, such as for a hard $\beta$-FPUT chain with $\beta>0$ as well as for a soft one with $\beta<0$.
Figure 2: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a heavy isotopic defect. The anti-localization, i.e., zeroing of the perturbation amplitude near $n=0$, is clearly observed for both linear and nonlinear systems with the defect and not observed for systems without the defect. The difference between nonlinear and linear solutions for the system with the defect is clearly recognizable everywhere excepting the anti-localization zone.
Figure 3: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain. (a) The solutions for a system with a heavy isotopic defect; (b) the solutions for the corresponding uniform system. In both sub-plots, the order of decay for the nonlinear system can be estimated as the same as it is observed in the corresponding linear system. Oscillation asymmetry is clearly recognizable for the nonlinear system with the defect.
Figure 4: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in an infinite $\beta$-FPUT chain and the corresponding linear chain with and without a light isotopic defect. The anti-localization in systems with the defect co-exists with impurity-induced localized mode. The latter forms a peak of the amplitude near $n=0$
Figure 5: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain and in the corresponding linear chain in the case of a light isotopic defect. For both linear and nonlinear systems, a non-vanishing localized oscillation is observed. The frequency shift observed for the nonlinear system confirms the nonlinear character of the oscillation
IV An infinite $\beta$-FPUT chain with a mass-spring inclusion
As we can see from the previous Sect. III, the anti-localization exists in linear and nonlinear systems with an isotopic defect. The anti-localization is observed for both cases $m>1$ and $m<1$, though, in the latter case, the anti-localization co-exists with the localization. The boundary in the system parameter space between these two cases corresponds to the uniform system, where there is no anti-localization. In previous study Shishkina et al. (2023), we have shown that besides the uniform system, the anti-localization disappears in linear non-uniform systems with parameters, which correspond to such a boundary separating domains in the system parameter space where the localization is observed and not observed. The linear uniform system is just a particular case of such a system with the boundary parameter values. In this section, we want to check if the anti-localization disappears for non-uniform nonlinear systems with such boundary values of their parameters. To achieve this, we consider the chain with parameters satisfying Eq. (12), equipped with a mass-spring inclusion for which Eq. (11) and
$K\neq 1$ (15)
are fulfilled. The defect is subjected to a pulse loading at the position defined by Eq. (14). In the corresponding linear system with $B=\beta=0$ the impurity-induced localization is observed if and only if the following inequality is fulfilled Montroll and Potts (1955); Yu (2019); Shishkina and Gavrilov (2023b):
$K>K_{0},$ (16)
$K_{0}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\frac{2m}{1+m}.$ (17)
Inequality (16) defines a domain in the problem parameter space, which we call the localization domain. Now consider a nonlinear system with a fixed value $B=\beta$, and fixed $m$ as we have taken in Sect. III. Now we try to determine the boundary value of the parameter $K$, which separates nonlinear systems where the impurity-induced localization is observed or not observed. The numeric experiment shows that such a value is very close to the value $K=K_{0}$ defined by Eq. (17) for the linear systems, see Figs. 68.
Figure 6: The particle velocity $v_{0}(t)$ versus time in the case of the mass-spring defect. Parameters of defect satisfy equation $K=K_{0}$. The order of decay for all three systems can be estimated as the same
Figure 7: The particle velocity $v_{n}(t)$ versus the spatial variable $n$ in the case of the mass-spring defect. Parameters of defect satisfy equation $K=K_{0}$. The anti-localization is absent in all three cases (the minimum at $n=0$ observed for the linear systems is not an amplitude minimum and related to the value of the phase $\phi t$)
Figure 8: The particle velocity $v_{0}(t)$ versus time in an infinite $\beta$-FPUT chain with a mass-spring defect. The orders of decay are clearly different in all three cases.
In Fig. 6 the numerical solution $v_{0}(t)$ versus the time $t$ for $\beta$-FPUT chain with a mass-spring defect is presented. The value of the linear spring stiffness is taken equal to the boundary value $K_{0}$. For reference, the same particle velocity in the corresponding linear system with the same defect, as well as the exact solution (9) for the uniform linear chain, are also presented  [^5] Unfortunately, we do not have here the reference asymptotic non-stationary solution for the corresponding linear system.. The order of decay for oscillation in both nonlinear systems with and without the defect can be estimated as $1/\sqrt{t}$ as it is observed according to Eqs. (35), (44) in the uniform linear chain. In Fig. 7, the particle velocities $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ are shown at a fixed value of time for the same three systems as in Fig. 6. The absence of the anti-localization in the system with the defect for which $K=K_{0}$ is clearly recognizable.
In Fig. 8 the same numerical solution $v_{0}(t)$ for $\beta$-FPUT chain with a mass-spring defect with $K=K_{0}$, which is already shown in Fig. 6, is compared with analogous solutions with $K=0.9K_{0}$ and $K=1.1K_{0}$. We can see that the solution corresponding to $K=0.9K_{0}$ vanishes faster than one taken for $K=K_{0}$. The corresponding order of decay is expected to be $1/\sqrt{t^{3}}$, see Sect. III. On the other hand, the solution corresponding to $K=1.1K_{0}$ is a non-vanishing one, i.e., an impurity induced localized oscillation is observed.
Thus, we have shown that the anti-localization is not observed in both linear and nonlinear systems with a mass-spring defect if the system parameters correspond to the boundary of the localization domain in the problem parameter space. On the other hand, the anti-localization is observed if the system parameters lie outside this boundary.
V A semi-infinite $\beta$-FPUT chain
Assuming that $K=0$, $B=0$, we split an infinite chain into two uncoupled semi-infinite chains. Thus, a semi-infinite chain loaded at the end $n=N=1$ is a particular case of a chain with a defect.
For the corresponding linear semi-infinite system, the effect that we call the anti-localization was first time discovered, as far as we know  [^6] We are grateful to S.D. Liazhkov., much earlier in Jackson (1978) and explained by specific peculiarities of linear systems. The basic formulas describing anti-localized wave-field in the linear case the reader can find in Appendix C. Let us check if the anti-localization exists in a semi-infinite $\beta$-FPUT chain.
The numerical solution, which corresponds to a nonlinear chain with $\beta\neq 0$ is presented in Fig. 9. For reference, the wave-field in the corresponding linear semi-infinite chain given by the exact solution (45) is also displayed. In Fig. 9a the particle velocities $v_{n}$ versus the spatial variable $n$ are presented. In Fig. 9b the particle velocity of the end particle $v_{1}(t)$ versus time is shown. One can see that the anti-localization is clearly visible for both linear and nonlinear systems.
Remark 1.
In Fig. 9a one can also observe a soliton-like wave propagating faster than the leading wave-front in the linear system. The solitons in a $\beta$-FPUT chain are discussed, e.g., in Vainchtein (2022).
Figure 9: The wave-field in the $\beta$-FPUT and linear chains loaded at the free end: (a) particle velocities $v_{n}$ versus the spatial variable $n$; (b) particle velocity of the end particle $v_{1}$ versus time. The anti-localization is clearly observed
VI A nonlinear spring in an infinite linear chain
In this section, we want to show that introducing the local nonlinearity into a linear system can also lead to a phenomenon similar to the anti-localization. We consider a uniform linear chain, i.e.,
$m=1,\quad K=1,\quad\beta=0.$ (18)
At first, consider the source at $N=0$ at the inclusion in the form of a hard nonlinear spring:
$B_{L}=0,\quad B_{R}>0.$ (19)
In Fig. 10 we present the spatial distribution of the particle velocities in the corresponding numeric experiment with pulse loading applied at the particle with number $n=N=0$ just to the left from the nonlinear spring. Sub-plots (a) & (b) correspond to the choices $B_{R}=4.0$ and $B_{R}=10.0$, respectively. For reference, the particle velocities in the corresponding linear system with $B_{L}=B_{R}=0$ given by Eq. (9) are also displayed in each sub-plot. One can see that the wave pattern observed in Fig. 10b is a bit similar to the one presented in Fig. 4, where the anti-localization co-exists with the impurity-induced localized oscillation. The clear difference is an expected asymmetry of the wave patterns in Fig. 10. In Fig. 10a we also observe weakening of the propagating component near the defect but not zeroing. Increasing the coefficient $B_{R}$ one can see that the amplitude of the propagating component near the defect decreases. Thus, we observe a weakening of the propagating component but, apparently, not an asymptotic one. The observed phenomenon can be called the incomplete anti-localization. Here we cannot estimate, even qualitatively, the order of decay for the anti-localized component due to the presence of the localized mode.
Figure 10: The particle velocity $v_{n}$ versus the spatial variable $n$ in an infinite linear chain with and without a nonlinear spring. The greater $B_{R}$, the tendency to the anti-localization is more clear
A possibility to eliminate the localized oscillation is to consider a pulse loading applied at some distance $N>0$ outside the defect. For such a case, in study Gavrilov and Shishkina (2024), we have shown that in the linear chain the anti-localization is observed in some one-sided neighborhood just behind the defect, which expands with time. The amplitude of the localized oscillation in the linear system exponentially decreases with $N$; therefore this oscillation is not observable for large enough $N$. The order of decay for the particle velocities in such a shadow zone is the same as it is observed when the loading is applied at a defect, i.e., $t^{-3/2}$. In Fig. 11, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ in an infinite linear chain, with an inclusion in the form of one strongly nonlinear spring, is shown for a fixed value of time. Here, we take an extremely large value $B_{R}=1000$ to obtain a clearly visible effect. For reference, the wave-field in the uniform linear chain given by Eq. (9), is also displayed. One can compare Fig. 11 with Fig. 1 in Shishkina and Gavrilov (2023a), where the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for the analogous problem for the linear chain with an isotopic defect. The plots are very similar, and the shadow zones behind the defect are clearly recognizable in both cases.
Figure 11: The particle velocity $v_{n}$ versus the spatial variable $n$ in an infinite linear chain with a hard, strongly nonlinear spring. The source position $n=N$ is indicated by the vertical magenta solid line. One can clearly see a shadow behind the defect at $n=0$
In Fig. 12 the numerically found solution for the particle velocity $v_{-1}(t)$ versus the time $t$ is shown for the same system. The choice n=1n=-1 corresponds to the first particle behind the defect, which is not connected to the defective spring. One can see that the estimate for the order of the decay should be taken as $t^{-1/2}$, not $t^{-3/2}$ as usual when the (complete) anti-localization is observed. Thus, we really observe a weakening of the oscillation behind the defect, but not an asymptotic one. The leading order term of the large time asymptotics for the oscillation amplitude is relatively small, but not zero.
Figure 12: The particle velocity $v_{-1}$ versus time in an infinite linear chain with a nonlinear spring. The order of decay can be estimated as $1/\sqrt{t}$
VII Conclusion
In the paper, we have presented the numerical results for the investigation of several strongly nonlinear problems concerning the quasi-waves propagation in an infinitely long $\beta$-FPUT chain with a defect. Though we have demonstrated the results, which correspond to a specific choice of the problem parameters, in our work, we have considered a wide band of them. This allows us to claim that the phenomenon of anti-localization of non-stationary waves is not somehow related to the linearity of the system under consideration. Anti-localization is natural for both linear and nonlinear systems with defects.
Depending on the parameters of the system, we have the following alternatives:
  1. 1.
    An anti-localized propagating component of a wave-field may co-exist with an impurity-induced localized wave; see, e.g., Fig. 4, which corresponds to a light isotopic defect. The localized mode forms a peak of the amplitude near a defect, but this peak is observed inside a zone, where the oscillation amplitude is almost zero.
  2. 2.
    An anti-localized propagating component can be observed near a defect in the absence of impurity-induced localized modes; see, e.g., Fig. 2, which corresponds to a heavy isotopic defect.
  3. 3.
    The anti-localization is not observed for systems with parameters at the boundary of the localization domain  [^7] A domain in the problem parameter space where the wave localization occurs.; see, e.g., Fig. 7, which corresponds to a mass-spring defect. Thus, we have shown that the absence of the anti-localization in a uniform system is not related to the uniformity itself. A uniform system is just the particular case of a system with such boundary parameters.
These results were shown previously in Shishkina et al. (2023) for a continuum linear system in an analytic way. In this paper, we have demonstrated numerically an analogous behavior of strongly nonlinear systems.
In Sect. VI, we have demonstrated that embedding of the point strong nonlinearities into a linear chain can also lead to a weakening of the propagating component of the wave-field in a neighborhood of such a defect. Though this effect is quite similar to one which we call the anti-localization, in the latter case, we observe weakening but not zeroing of the leading-order term. This can be called ‘‘the incomplete anti-localization’’.
Although the conclusions in the current paper are very similar to ones previously obtained for the linear case, we claim that they cannot be firmly deduced from any linear consideration. The observations that support this point of view are as follows:
  • The considered nonlinearity clearly quantitatively alters the solution at least everywhere excepting the neighborhood of the defect; see, e.g., Figs. 2, 9a;
  • The considered nonlinearity clearly quantitatively alters solution at the defect for the finite (not very small and not very large) values of time; see, e.g., Figs. 3a, 9b;
  • In the case of a heavy isotopic defect, nonlinearity qualitatively alters the solution at the defect for finite and large times. Namely, oscillation asymmetry is observed; see Fig. 3a. This effect is impossible Shishkina and Gavrilov (2023a) in the linear system.
  • Anti-localization can be observed in a neighborhood of a nonlinear impurity-induced localized mode; see, e.g., Fig. 4;
  • The phenomenon of the incomplete anti-localization discussed in Sect. VI is observed for a nonlinear system, which is uniform in the linear approximation;
  • The anti-localization in a linear semi-infinite chain has been demonstrated in the strongly nonlinear case; see Fig. 9; although before (Jackson (1978), p. 137) this effect was explained as a specific peculiarity of a linear system;
In our opinion, a nonlinearity, of course, could destroy the anti-localization as suggested in Jackson (1978). Indeed, in domains where localized waves are not observed, the linear equations are applicable for large times. However, the initial and boundary conditions for these equations are completely different from ones that we consider investigating the anti-localization in linear systems. Nonlinear localized waves of a finite amplitude (impurity-induced ones and solitary waves at the leading wave-fronts), in principle, could destroy the anti-localization. These localized structures are the sources, from which the perturbations propagate into a neighborhood of the defect. In Gavrilov and Shishkina (2024) we have shown that in the linear case a source outside the defect destroys anti-localization inside a zone between the source and the defect. The essentially nonlinear stage of the motion, which corresponds to finite values of time, also can affect defect’s oscillation at large times, see Fig. 3a. Nevertheless, the investigation in this paper shows that the anti-localization survives in nonlinear systems. Thus, the anti-localization can also be observed in weakly nonlinear systems for very large times, i.e., in the thermodynamic limit.
Note that introducing nonlinearity into a linear system possessing a localized mode with the corresponding frequency inside a stop-band, in some cases, can lead to destroying the mode. This happens at least in the case when the stop-band correspond to smaller values of frequency compared with the pass-band Kaplunov and Nolde (2008); Indeitsev and Osipova (2004) due to the emergence of super-harmonics, whose frequencies are multiples of the localized mode frequency. These frequencies can leave the stop-band yielding the leakage of energy of the localized mode. In this paper, we consider systems with a stop-band corresponding to higher values of the frequency compared with a pass-band. This situation is quite usual for discrete systems with zero on-site potential. The anti-localization is the phenomenon, which is closely related to the localization; thus, in nonlinear systems where a stop-band corresponds to small values of the frequency (this is typical for the continuous case) the anti-localization requires an additional investigation.
Acknowledgements
The authors are grateful to E.F. Grekova, S.D. Liazhkov, Yu.A. Mochalova, A.V. Porubov for useful and stimulating discussions. The research is supported by the Ministry of Science and Higher Education of the Russian Federation (project 124040800009-8).
Appendix A Non-dimensionalization
In the dimensional form, the equations of motion are:
$m_{\infty}\frac{d^{2}{\tilde{u}}_{n}}{d\tilde{t}^{2}}-K_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})-(\tilde{u}_{n}-\tilde{u}_{n-1})\big{)}\\ -B_{\infty}\big{(}(\tilde{u}_{n+1}-\tilde{u}_{n})^{3}-(\tilde{u}_{n}-\tilde{u}_{n-1})^{3}\big{)}\\ =\Big{(}(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})-(\tilde{u}_{0}-\tilde{u}_{-1})\big{)}\\ +\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\big{)}\Big{)}\delta_{n}\\ -(\tilde{K}-K_{\infty})\big{(}(\tilde{u}_{1}-\tilde{u}_{0})\delta_{n-1}-(\tilde{u}_{0}-\tilde{u}_{-1})\delta_{n+1}\big{)}\\ -\big{(}(\tilde{B}_{R}-B_{\infty})(\tilde{u}_{1}-\tilde{u}_{0})^{3}\delta_{n-1}\\ {-(\tilde{B}_{L}-B_{\infty})(\tilde{u}_{0}-\tilde{u}_{-1})^{3}\delta_{n+1}\big{)}}\\ -(\tilde{m}-m_{\infty})\frac{d^{2}{\tilde{u}}_{0}}{d\tilde{t}^{2}}\delta_{n}.$ (20)
[…]
The initial conditions are
$\tilde{u}_{n}(0)=0,\quad\frac{d\tilde{u}_{n}}{d\tilde{t}}(0)=\frac{\tilde{p}\delta_{n-N}}{m_{\infty}+\delta_{n}(\tilde{m}-m_{\infty})}.$ (21)
Initial conditions in the form of Eq. (21) correspond to the pulse loading $\tilde{p}\delta(t)$ applied to the particle with label $N$. Here $\tilde{p}$ is the magnitude of the loading, $\delta(t)$ is the Dirac delta-function.
Put
$\begin{gathered}\omega=\sqrt{\frac{K_{\infty}}{m_{\infty}}},\quad t=\omega\tilde{t},\quad u_{n}=\frac{\tilde{u}_{n}K_{\infty}}{\tilde{p}\omega},\\ K=\frac{\tilde{K}}{K_{\infty}},\quad m=\frac{\tilde{m}}{m_{\infty}},\\ B_{L}=\frac{\tilde{B}_{L}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad B_{R}=\frac{\tilde{B}_{R}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}},\quad\beta=\frac{B_{\infty}\tilde{p}^{2}}{m_{\infty}K_{\infty}^{2}}.\end{gathered}$ (22)
Using dimensionless quantities (22), Eq. (20) may be rewritten in the following dimensionless form:
$\ddot{u}_{n}-(u_{n+1}-2u_{n}+u_{n-1})\\ -\beta\big{(}(u_{n+1}-u_{n})^{3}-(u_{n}-u_{n-1})^{3}\big{)}\\ =\big{(}(K-1)\big{(}(u_{1}-u_{0})-(u_{0}-u_{-1})\big{)}\\ +\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\big{)}\big{)}\delta_{n}\\ -(K-1)\big{(}(u_{1}-u_{0})\delta_{n-1}-(u_{0}-u_{-1})\delta_{n+1}\big{)}\\ -\big{(}(B_{R}-\beta)(u_{1}-u_{0})^{3}\delta_{n-1}-(B_{L}-\beta)(u_{0}-u_{-1})^{3}\delta_{n+1}\big{)}\\ -(m-1)\ddot{u}_{0}\delta_{n},$ (23)
which is equivalent to Eqs. (1)–(4) if notation (5) is accepted. The initial conditions (21) can be reformulated in the form of Eq. (6).
Appendix B An infinite linear chain with an isotopic defect
For the reader’s convenience, we provide here some important formulas obtained in Shishkina and Gavrilov (2023a) that describes the solution of the problem considered in Sect. III, Eqs. (10)–(14) in the linear case $\beta=0$. The exact solution for the particle velocities in the integral form is an inverse Fourier transform:
${v}_{n}=-\frac{\mathrm{i}}{2\pi}\left(\int_{\mathbb{P}}+\int_{\mathbb{S}}\right)\Omega\mathcal{G}_{n}(\Omega)\mathrm{e}^{-\mathrm{i}\Omega t}\,d\Omega+\mathrm{c.c.}\\ =I_{n}^{\mathrm{pass}}+I_{n}^{\mathrm{stop}}+\mathrm{c.c.}=v_{n}^{\mathrm{pass}}+v_{n}^{\mathrm{stop}},$ (24)
where $\mathcal{G}_{n}(\Omega)$ is the corresponding Green function in the frequency domain:
$\mathcal{G}_{n}(\Omega)=-\frac{\mathrm{e}^{\mathrm{i}|n|\operatorname{sign}\Omega\,\arccos\frac{2-\Omega^{2}}{2}}}{(m-1)\Omega^{2}+\mathrm{i}\Omega\sqrt{4-\Omega^{2}}},\ \ \Omega\in\mathbb{P};$ (25)
$\mathcal{G}_{n}(\Omega)=\frac{(-1)^{|n|}2^{|n|}}{\Phi^{|n|-1}(\Omega)\big{(}(-m\Omega^{2}+2)\Phi(\Omega)+4\big{)}},\ \ \Omega\in\mathbb{S};$ (26)
$\mathrm{c.c.}$ are the complex conjugate terms. Here,
$\Phi(\Omega)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\Omega^{2}-2+|\Omega|\sqrt{\Omega^{2}-4};$ (27)
$\mathbb{P}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}[0,\Omega_{\ast}],\qquad\mathbb{S}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}(\Omega_{\ast},\infty)$ (28)
are the pass-band and the stop-band of the whole non-negative frequency band $\Omega\in[0,+\infty]$; $\Omega_{\ast}=2$ is the cut-off frequency which separates the bands.
Remark 2.
In Shishkina and Gavrilov (2023a) mostly the heat, i.e., the kinetic energy, transfer processes are under consideration. Thus, the solution is obtained for the particle velocities, not for the displacements. This is also useful because the displacements involve a non-vanishing non-oscillating component (a rigid body mode of motion).
The integral $I_{n}^{\mathrm{stop}}$ describes, in particular, an impurity-induced non-vanishing localized oscillation, which exists only in the case of a light defect
$0<m<1.$ (29)
If the last inequality is true, in the interval $(0,\Omega_{\ast})$ there exists a simple root
$\Omega_{0}=\frac{2}{\sqrt{m(2-m)}}$ (30)
of the denominator for $\mathcal{G}^{\mathrm{stop}}$. Therefore, integral $I^{\mathrm{stop}}$ does not exist in the classical sense. In the latter case, integral $I^{\mathrm{stop}}$ should be considered as the Fourier transform for a generalized function. To regularize this, we can apply Shishkina et al. (2023); Shishkina and Gavrilov (2023a) the limit absorption principle. Finally, for $t\to\infty$ at a fixed position $n$ one gets
$v_{n}^{\mathrm{stop}}=I^{\mathrm{stop}}+\mathrm{c.c.}=H(1-m)L_{n}(t)+O(t^{-1}),$ (31)
${L}_{n}=\frac{2(m-1)(-1)^{|n|+1}m^{|n|-1}H(1-m)}{(2-m)^{|n|+1}}\cos\Omega_{0}t,$ (32)
where $t\to\infty$, see Shishkina and Gavrilov (2023a). Here $L_{n}(t)$ is the localized non-vanishing oscillation, $H(\cdot)$ is the Heaviside step-function.
The phenomenon of the anti-localization is mostly related to the integral $I^{\mathrm{pass}}$, which describes the propagating part of the wave-field. Following to Gavrilov (2022), we have estimated it in Shishkina and Gavrilov (2023a) on a moving at an arbitrary sub-critical speed $w$ point of observation. Taking into account that solution (24) clearly is an even function of $n$, put
$|n|=wt,$ (33)
$0<w<1$ (34)
in the expression for $I^{\mathrm{pass}}$ defined by Eq. (24). The obtained integral can be estimated for $t\to\infty$ using the method of stationary phase. This provides an asymptotic continuum solution:
$v_{n}^{\mathrm{pass}}=I^{\mathrm{pass}}+\mathrm{c.c.}\\ =\frac{{A}(w)}{\sqrt{t}}\cos\Big{(}\phi(w)\,t+\frac{\pi}{4}+\psi(w)\Big{)}+O(t^{-3/2}),$ (35)
[…]
$\phi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}2(w\arccos w-\sqrt{1-w^{2}}),$ (40)
$\psi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\arctan\frac{(m-1)\sqrt{1-w^{2}}}{w}.$ (41)
Note that for $w>1$ there is no stationary point, and the corresponding term of order $1/\sqrt{t}$ is zero. This fact is taken into account by introducing the multiplier H(1w)H(1-w) in the numerator of the right-hand side of Eq. (39). In any system with an isotopic impurity, i.e., in the case (11), we have:
${A}(w)=\frac{\sqrt{2}w}{\sqrt{\pi}|m-1|}+O(w^{3}),\quad w\to 0;\qquad A(0)=0;$ (42)
and, thus, the amplitude $A(w)$ of the propagating part $v^{\mathrm{pass}}$ for the particle velocities is small in a certain expanding (since $w=|x|/t$) neighborhood of the impurity. We call this phenomenon the anti-localization of non-stationary waves. Finally, the asymptotic expansion for the right-hand side of Eq. (24) at $n=0$ (just at the impurity) has the following form
$v_{0}(t)=H(M-K)L_{0}(t)+\frac{\cos\left(2t-\frac{3\pi}{4}\right)}{2\sqrt{\pi}(m-1)^{2}\,t^{3/2}}\\ +o(t^{-3/2}).$ (43)
The first term in the right-hand side of Eq. (43) is the contribution from the poles $\pm\Omega_{0}$ and describes localized oscillation, which exists if and only if $m<1$. The second term is the anti-localized part of the wave-field at the inclusion expressed as the total contribution from the cut-off frequency $\Omega_{\ast}$ for both integrals $I^{\mathrm{stop}}$ and $I^{\mathrm{pass}}$.
In the absence of the impurity, the anti-localization is not observed:
$A(w)=\frac{1}{\sqrt{\pi{\sqrt{1-w^{2}}}}}=\frac{1}{\sqrt{\pi}}+O(w^{2}),\quad w\to 0;\\ \qquad A(0)\neq 0.$ (44)
Though the last formula can be formally derived from Eq. (39), this way is incorrect since Eq. (39) is obtained assuming Eq. (11). The correct way to obtain formula (44) is to use the exact solution (9) Gavrilov (2022). Thus, the anti-localization in the linear system is an asymptotic weakening of the propagating component of the wave-field. The order of decay for the vanishing component is $1/\sqrt{t}$ in the uniform system and $1/\sqrt{t^{3}}$ in the system with a defect (near the defect).
Appendix C A semi-infinite linear chain
The exact solution is Liazhkov (2025)
$v_{n}=J_{2(n-1)}(2t)+J_{2n}(2t),\qquad n\in\mathbb{N}.$ (45)
The anti-localization in the semi-infinite linear chain ($\beta=0$) was demonstrated in study Gavrilov et al. (2023), where the asymptotic solution for $v_{n}$ in the form of the right-hand side of Eq. (35) wherein
${A}(w)=\frac{2w}{\sqrt{\pi}\sqrt[4]{1-w^{2}}},$ (46)
$\psi=\arctan\frac{\sqrt{1-w^{2}}}{w},$ (47)
$w=\frac{|n-1|}{t}$ (48)
was obtained. The expression for $\phi$ coincides with (40). Clearly, $A(0)=0$, which implies the anti-localization effect. For the end particle, one has the anti-localized solution Gavrilov et al. (2023):
$v_{1}=\frac{\sin(2t-\frac{\pi}{4})}{\sqrt{\pi}t^{3/2}}+o(t^{-3/2}).$ (49)